OpenStudy (anonymous):

The volume of a cube is increasing at a rate of 10cm^3/min. How fast is the surface area increasing when the length of an edge is 30cm?

OpenStudy (jdoe0001):

\( \cfrac{dV}{dt}=+10\qquad \qquad V=lwh\qquad \begin{cases} l=x\\ w=x\\ h=x \end{cases}\implies V=x\cdot x\cdot x\implies V=x^3\) in a cube, all sides are equal, thus length, height and width are all "x", thus you'd end up with something like \(x^3\)

OpenStudy (anonymous):

let x be the edge of the cube volume \[V=x^3\] \[\frac{ dV }{ dt }=3x^2\frac{ dx }{ dt },\] when x=30 cm \[10=3*30^2\frac{ dx }{ dt }\] \[\frac{ dx }{ dt }=?\] surface area \[S=6x^2\] \[\frac{ dS }{ dt }=6*2x \frac{ dx }{ dt }\] when x=30 cm find \[\frac{ dS }{ dt }\]

OpenStudy (jdoe0001):

yeap, notice above as surjithayer said  the cube is just 6 squares, each of x * x, so the surface area of it will be \(\bf s = 6x^2\)